# Joint Self-Supervised Video Alignment and Action Segmentation

### v2

@@ 5 Conclusion
This paper presents a novel approach for joint self-supervised video alignment and action segmentation. We first develop a fused Gromov-Wasserstein optimal transport with a structural prior for self-supervised video alignment, outperforming prior works. Our single-task method trains efficiently on GPUs and needs few iterations to derive the optimal transport solution. Next, we extend our approach to a unified optimal transport framework for joint self-supervised video alignment and action segmentation, yielding similar video alignment yet better action segmentation results than prior works. Our multi-task method requires training and storing a single model and saves both time and memory usage. To our best knowledge, our work is the first to explore the relationship between video alignment and action segmentation. Our future works will explore deep supervision [41, 42], complex weighting in multi-task learning [29, 9], and other potential applications (e.g., joint keypoint matching and clustering [20, 56]).
@@ A.1 Implementation Details
For fair comparisons with previous self-supervised video alignment methods [19, 26, 45, 17], our VAOT and VASOT approaches for video alignment utilize a ResNet-50 encoder network. Please refer to Dwibedi et al. [19] for additional details on the encode network. We provide the hyperparameter settings of our VAOT and VASOT approaches for video alignment in Tab. 4. Note that our VASOT approach for video alignment includes both VAOT and ASOT [80] modules, which share the same hyperparameter settings shown in the VASOT column in Tab. 4.
In addition, following state-of-the-art self-supervised action segmentation methods [34, 35, 80], our VASOT approach for action segmentation employs a 2-layer MLP encoder network for fair comparison purposes. Please see Kukleva et al. [34] for more details on the encoder network. The hyperparameter settings of our VASOT approach for action segmentation are included in Tab. 5. Note that both VAOT and ASOT [80] components in our VASOT approach for action segmentation have the same hyperparameter settings presented in the VASOT column in Tab. 5.
@@ A.4 Multi-Action Video Alignment Results
Penn Action SAL [48] 68.15 39.03 47.44
TCN [57] 68.09 38.34 54.17
TCC [19] 74.39 59.14 64.08
LAV [26] 78.68 62.52 68.35
VAVA [45] 80.25 64.82 76.20
GTCC [17] 73.90 69.70 60.70
@@ A.5 Per-Video Action Segmentation Results
Per-Video TWF [55] 62.7 / 49.8 / 42.3 56.7 / 48.2 / - 66.8 / 56.4 / 48.7 71.7 / - / - 73.3 / 67.7 / 57.7
ABD [18] 64.0 / 52.3 / - 67.2 / 49.2 / - 71.8 / - / - 71.22/ - / - - / - / -
ASOT [80] 63.3 / 53.5 / 35.9 71.2 / 63.3 / 47.8 64.3 / 51.1 / 33.4 64.5 / 58.9 / 33.0 73.0 / 68.4 / 47.6
@@
We perform Hungarian matching over the entire dataset to obtain the full-dataset action segmentation results in Tab. 3 of the main paper. In the following, we benchmark our VASOT approach against previous unsupervised action segmentation methods [55, 18] which conduct Hungarian matching and evaluate per video. Per-video matching and evaluation do not require clusters across all videos and hence tend to yield better results. Tab. 8 presents the results in the per-video matching and evaluation setting. It can be seen from the results that our VASOT approach achieves the best overall performance in this experiment setting. Especially, on F1 score, our VASOT approach consistently performs the best across all datasets.
@@ A.6 Qualitative Results
We provide some qualitative results in this section. Firstly, Fig. 7 presents the frame retrieval results of our VAOT approach and LAV [26] on Penn Action. From Fig. 7, our VAOT approach retrieves all 5 correct frames with the same action (Bat swung back fully) as the query image, whereas LAV [26] retrieves all incorrect frames (Bat hits ball), highlighted by red ovals.
Secondly, Fig. 8 illustrates the t-SNE visualizations of the learned frame embeddings of two Pouring videos by our VAOT approach and LAV [26]. In Fig. 8(a), the embeddings by LAV [26] form locally continuous but globally fragmented trajectories, with visible gaps in earlier and later frames. In contrast, for our VAOT approach in Fig. 8(b), the embeddings of corresponding frames from both videos are spatially closer and follow smoother trajectories. This suggests that our VAOT approach can effectively capture both local and global temporal information, resulting in more temporally consistent embeddings.
Lastly, Fig. 9 shows the action segmentation results by our VASOT approach and ASOT [80] on a Desktop Assembly video and a YouTube Instructions video. The Desktop Assembly dataset is relatively balanced, with actions of more uniform durations, whereas the YouTube Instructions dataset is more unbalanced, with some actions being significantly longer or shorter than others. From Fig. 9, our VASOT approach can effectively handle both cases, yielding segmentations which align more closely with ground truth than ASOT [80].

### v3

@@ 5 Conclusion
This paper presents a novel approach for joint self-supervised video alignment and action segmentation. We first develop a fused Gromov-Wasserstein optimal transport with a structural prior for self-supervised video alignment, outperforming prior works. Our single-task method trains efficiently on GPUs and needs few iterations to derive the optimal transport solution. Next, we extend our approach to a unified optimal transport framework for joint self-supervised video alignment and action segmentation, yielding similar video alignment yet better action segmentation results than prior works. Our multi-task method requires training and storing a single model and saves both time and memory usage. To our best knowledge, our work is the first to explore the relationship between video alignment and action segmentation. Our future works will explore deep supervision [li2017deep, li2018deep], complex weighting in multi-task learning [kendall2018multi, chen2018gradnorm], and other potential applications (e.g., joint keypoint matching and clustering [estrada2009appearance, sarlin2020superglue]).
@@ A.1 Implementation Details
For fair comparisons with previous self-supervised video alignment methods [dwibedi2019tcc, haresh2021learning, liu2022learning, donahue2024learning], our VAOT and VASOT approaches for video alignment utilize a ResNet-50 encoder network. Please refer to Dwibedi et al. [dwibedi2019tcc] for additional details on the encode network. We provide the hyperparameter settings of our VAOT and VASOT approaches for video alignment in Tab. 4. Note that our VASOT approach for video alignment includes both VAOT and ASOT [xu2024temporally] modules, which share the same hyperparameter settings shown in the VASOT column in Tab. 4.
In addition, following state-of-the-art self-supervised action segmentation methods [kukleva2019unsupervised, kumar2022unsupervised, xu2024temporally], our VASOT approach for action segmentation employs a 2-layer MLP encoder network for fair comparison purposes. Please see Kukleva et al. [kukleva2019unsupervised] for more details on the encoder network. The hyperparameter settings of our VASOT approach for action segmentation are included in Tab. 5. Note that both VAOT and ASOT [xu2024temporally] components in our VASOT approach for action segmentation have the same hyperparameter settings presented in the VASOT column in Tab. 5.
@@ A.4 Multi-Action Video Alignment Results
Penn Action SAL [misra2016shuffle] 68.15 39.03 47.44
TCN [sermanet2018time] 68.09 38.34 54.17
TCC [dwibedi2019tcc] 74.39 59.14 64.08
LAV [haresh2021learning] 78.68 62.52 68.35
VAVA [liu2022learning] 80.25 64.82 76.20
GTCC [donahue2024learning] 73.90 69.70 60.70
@@ A.5 Per-Video Action Segmentation Results
Per-Video TWF [sarfraz2021temporally] 62.7 / 49.8 / 42.3 56.7 / 48.2 / - 66.8 / 56.4 / 48.7 71.7 / - / - 73.3 / 67.7 / 57.7
ABD [du2022fast] 64.0 / 52.3 / - 67.2 / 49.2 / - 71.8 / - / - 71.22/ - / - - / - / -
ASOT [xu2024temporally] 63.3 / 53.5 / 35.9 71.2 / 63.3 / 47.8 64.3 / 51.1 / 33.4 64.5 / 58.9 / 33.0 73.0 / 68.4 / 47.6
@@
We perform Hungarian matching over the entire dataset to obtain the full-dataset action segmentation results in Tab. 3 of the main paper. In the following, we benchmark our VASOT approach against previous unsupervised action segmentation methods [sarfraz2021temporally, du2022fast] which conduct Hungarian matching and evaluate per video. Per-video matching and evaluation do not require clusters across all videos and hence tend to yield better results. Tab. 8 presents the results in the per-video matching and evaluation setting. It can be seen from the results that our VASOT approach achieves the best overall performance in this experiment setting. Especially, on F1 score, our VASOT approach consistently performs the best across all datasets.
@@ A.6 Qualitative Results
We provide some qualitative results in this section. Firstly, Fig. 7 presents the frame retrieval results of our VAOT approach and LAV [haresh2021learning] on Penn Action. From Fig. 7, our VAOT approach retrieves all 5 correct frames with the same action (Bat swung back fully) as the query image, whereas LAV [haresh2021learning] retrieves all incorrect frames (Bat hits ball), highlighted by red ovals.
Secondly, Fig. 8 illustrates the t-SNE visualizations of the learned frame embeddings of two Pouring videos by our VAOT approach and LAV [haresh2021learning]. In Fig. 8(a), the embeddings by LAV [haresh2021learning] form locally continuous but globally fragmented trajectories, with visible gaps in earlier and later frames. In contrast, for our VAOT approach in Fig. 8(b), the embeddings of corresponding frames from both videos are spatially closer and follow smoother trajectories. This suggests that our VAOT approach can effectively capture both local and global temporal information, resulting in more temporally consistent embeddings.
Lastly, Fig. 9 shows the action segmentation results by our VASOT approach and ASOT [xu2024temporally] on a Desktop Assembly video and a YouTube Instructions video. The Desktop Assembly dataset is relatively balanced, with actions of more uniform durations, whereas the YouTube Instructions dataset is more unbalanced, with some actions being significantly longer or shorter than others. From Fig. 9, our VASOT approach can effectively handle both cases, yielding segmentations which align more closely with ground truth than ASOT [xu2024temporally].